OpenStudy (anonymous):

2 sqrt(x + 2 − 3 = 7)

OpenStudy (anonymous):

\[2\sqrt{x+2-3}=7\]

OpenStudy (anonymous):

sqrt(x+2-3)=7/2 take squqre on both sides x+2-3=49/4 x=(49/4)+3-2